# artificial intelligence CLI command-line interface

### v2

@@ 2.1 Watermarking and Steganography
Table 1: Digital Watermarking Definitions
@@ 2.2 Notation and Unicode Whitespace Characters
Table 2: Whitespace Evaluation Based on [28]
@@
Table 3: Overview of Notation, Following [27]
@@
$CT$ Cover text
$SM$ Secret message text
@@
$CT_{SM}$ Text containing a hidden secret message
@@ 3.1.8 CovertSYS
Table 4: Overview of Related Methods
@@
Innamark [19, 45] 2025 Paper, Software Replaces whitespace with a specific set of robust, similar-looking whitespace characters.
@@ 4 Proposed Method
In this section, we present Innamark, our invisible and blind information-hiding technique for plain text. Since existing methods either lack robustness in some applications, increase the number of characters, or are recognizable by humans, they are unsuitable for the LLM use case described in Section 1. Therefore, this section presents the embedding and extraction method in Algorithm 1 and Algorithm 2 based on our nomenclature introduced in Section 2.2. It further includes a concrete example based on some Lorem ipsum dummy text and information about our implemented prototypes.
@@ 4.1 Embedding
1:  $\triangleright$ Insert tag
42:  $SM_{bytes},SM_{H}\leftarrow applyTag(SM_{bytes},\theta)$
3:  $\triangleright$ Encode secret message
4:  $d\leftarrow\left\lceil\frac{\log_{2}2^{8}}{\log_{2}|\mathcal{A}_{-}|}\right\rceil$
@@
9:        $SM_{H}\leftarrow SM_{H}+a_{r+1}$triangleright $a_{r+1}\in\mathcal{A}_{-}$
@@
16:        if $i\leq|SM_{H}|$ then
17:           $w_{H_{i}}\leftarrow SM_{H_{i}}$triangleright $w_{H_{i}}\in\mathcal{A}_{-}$
18:           $CT_{SM}\leftarrow CT_{SM}+w_{H_{i}}$
19:           $i\leftarrow i+1$
20:        else
21:           $CT_{SM}\leftarrow CT_{SM}+\phi$
22:           $i\leftarrow 1$
23:        end if
24:     else
25:        $CT_{SM}\leftarrow CT_{SM}+c$
26:     end if
27:  end for
28:  return  $CT_{SM}$
Algorithm 1 Embedding, $Emb(CT,SM_{bytes},\theta)$
The embedding function $Emb(CT,SM_{bytes},\theta)$ begins by analyzing the configuration parameter $\theta$. This specifies the type of the secret message, which we call a InnamarkTag. It offers optional functionalities that can be enabled in any combination by the end user, namely, encryption, compression, hashing, and error-correcting codes. Depending on the user’s choice, $\theta$ defines the type used to calculate the tag. The tag has a fixed length of one byte and is returned by the $applyTag$ method with the updated secret message, depending on the user’s choice. Each bit in the tag indicates whether an option, such as compression, is enabled ($1$) or not (00). Thus, the tag describes the format of the secret message, similar to tags or headers in network packets [46]. Fig.  1 provides an overview of the InnamarkTag structure, the definition of each bit in the tag, and an example of a tag with enabled compression (second compression bit set to one) and error correction (fourth CRC32 bit set to one).
@@
An InnamarkTag starts with the separator character $\phi$, followed by the tag with a fixed length of one byte, followed by an optional prefix whose content depends on the tag, followed by the secret message itself.
Next, the encoding process starts transforming the secret message with the InnamarkTag tag and optional prefixes $SM_{bytes}$ into a sequence of whitespace characters $SM_{H}$ from our homoglyph alphabet $\mathcal{A}_{-}$. Since $|\mathcal{A}_{-}|=4$, where ||$\cdot$|| ⋅ | denotes the cardinality, each byte of the secret message is represented by four elements of $\mathcal{A}_{-}$ because
@@
The final version of the cover text with the hidden secret message $CT_{SM}$ is created by replacing all $\delta$ successively with the elements of $SM_{H}$. If, on the one hand, the input cover text $CT$ does not have any normal whitespace $\delta$, e.g., if the algorithm has already been applied to it, it is not possible to embed the secret message. If, on the other hand, $CT$ has more $\delta$ characters than $|SM_{H}|$, the next $\delta$ is replaced with the separator character $\phi$, and the insertion process starts again until all $\delta$ are replaced to include the secret message multiple times. Thus, the resulting text does not contain standard space characters:
@@
An example of the proposed embedding algorithm is shown in Fig. 2. The secret message $SM=$ “John” is to be hidden inside the cover text $CT=$ “Lorem ipsum …” with an empty configuration parameter $\theta$ to use the default tag without compression, hashing, or error correction. Each character of $SM$ is encoded into its byte representation $SM_{bytes}$ and transformed into the whitespace alphabet. In this case, the first letter “J” of the secret message is represented by the UTF-8 hexadecimal value 4A (U+004A), which equals the decimal value 74. Afterward, each value of $SM_{bytes}$ is encoded into $SM_{H}$ by transforming it into the alphabet $\mathcal{A}_{-}$ with a loop-based modulo operation, as described in Algorithm 1. This uses the length of the alphabet without separator characters as the divisor $d=|\mathcal{A}_{-}|=4$, as illustrated in Fig. 3. The remainder of each division operation indicates the index of the whitespace in $\mathcal{A}_{-}$ needed to build the complete sequence of whitespace characters $SM_{H}$ as a representation of the secret message $SM$. Next, each whitespace character $\delta$ of $CT$ is replaced with the corresponding space in $SM_{H}$. Since $CT$ has more whitespace characters than needed, the separator character $\phi$ is added, represented by the black U+2004 in Fig. 2. The insertion process starts again, and the algorithm inserts the identification tag and the first character “J” of the secret message a second time and stops after all whitespace characters are replaced.
@@ 4.2 Extraction
The third part decodes the hidden secret message $SM_{H}$ into its bit representation $SM_{bytes}$. The step size for the decoding part depends on the length of the alphabet without the separator character and is defined as
@@
10:  Extracted secret message bits ($SM_{bytes}$) or error
1:  $\triangleright$ Extract secret message
@@
11:  $\triangleright$ Analyze tag and prefix
12:  $SM_{H},error\leftarrow analyzeTag(SM_{H})$
@@
16:  $\triangleright$ Decode watermark
17:  $d\leftarrow\left\lceil\frac{\log_{2}2^{8}}{\log_{2}|\mathcal{A}_{-}|}\right\rceil$
@@
19:     for $y\leftarrow 0$ to $d-1$ do
20:        $a_{k}\leftarrow SM_{H_{i+y+1}}$triangleright $a_{k}\in\mathcal{A}_{-}$
21:        $b\leftarrow b+(k-1)\cdot d^{y}$triangleright $k\in[1,\dots,d]$
@@ 5.2 Capacity
[…]
@@
We assign all unbounded-capacity algorithms a value of $1.0$ for illustrate the unprescribed limit. Comparing the others, we find that Shiu et al. has the lowest embedding capacity, with $26/2514\approx 0.01$ bits/character, and that UniSpaCh has the highest of $1983/2514\approx 0.79$ bits/character. The other four bounded-capacity algorithms are close to each other and have an embedding capacity of around and below 0.1 bits/character, whereas our proposed Innamark method has a capacity of $93/2514\approx 0.04$ bits/character.
@@ 5.3 Imperceptibility
Imperceptibility or invisibility is the ability for a secret message to be concealed in a cover document without causing any visible abnormalities [21]. Therefore, this benchmark is more relevant for steganography use cases than watermarking [6, 11]. It is essential to distinguish between imperceptibility to humans and to machines, which can detect differences, for example, by using statistical metrics. Due to the varying perceptions of humans, this benchmark “is the most subjective of all the metrics” [34, p. 121]. In the following, we use four different metrics to measure the imperceptibility of the algorithms:
@@ 5.3.1 Jaro–Winkler Similarity
A standard numerical measurement used to compare the similarity between two character sequences is the Jaro–Winkler similarity, also known by the misleading name “Jaro–Winkler distance” [51]. It is often used to evaluate information-hiding techniques (see [42, 31, 27, 52, 53, 54]). The benchmark is based on the Jaro string comparator $\Phi$, for which a value of $1$ indicates that two strings $s_{1}$ and $s_{2}$ are identical and 00 indicates that the strings have no common characters [55] :
@@ 5.3.3 File Size
Humans and computer systems can detect a text file with a hidden secret message by its increased file size. In particular, suspicions may be raised if a file with a small amount of text has a large file size. Thus, Majeed et al. [8] argue for developing data-hiding methods that create results with minimal file sizes.
In our testbed, we compared the differences in file size between the original cover text and the text including a secret message. Although these benchmark criteria are closely related to the previous analysis of the number of characters, the results differ. This is related to a profound difference in the characters’ storage requirements. For example, Shazzad-Ur-Rahman et al.’s [43, 44] method replaces a classical small Latin letter “g” (U+0067) with the visually similar-looking mathematical alphanumeric symbol U+1D5C0 as part of the letterlike symbols in the Unicode standard [20]. Whereas the first needs one byte of storage space in the UTF-8 encoding, the second needs four bytes. Fig. 8 illustrates the mean absolute differences based on the two runs on our 1 000 000-article dataset. The negative value for short secret messages from the Shiu et al. algorithm indicates a decreased file size. This can be attributed to replacing whitespace characters with newlines, as seen in the results for the number of characters. Our proposed Innamark technique is just in the lower third in direct comparison because it is one of the few algorithms that embeds the secret message multiple times in the cover text for increased robustness.
@@ 5.3.4 Caret Navigation
In our testbed, we manually checked various texts with secret messages. We moved the caret from the start to the end to identify any unusual behavior. In cases of hanging caret positions or identifying unexpected characters, we classified the algorithm as visible because the caret attack revealed the secret message’s position. Table 5 shows an overview of the results with a justification for each algorithm as to why it is noticeable by users or not. It was found that only three algorithms (Rizzo et al., Lookalikes, and Innamark) are robust against caret navigations.
Table 5: Caret Navigation Attack Evaluation
@@ 5.4.1 Modification Robustness
Modifications to a file that may compromise the secret message, such as text insertions or deletions, are also known as tampering attacks in the information-hiding literature [33]. We have analyzed replacements, each consisting of a deletion followed by an insertion, for all implemented algorithms in batch runs for both the short and long secret messages on the 1 000 000-article dataset. For each article, we embedded the secret message and then replaced a block of 10% of the length of the original cover text with just one letter. The starting position of the replacement was randomly chosen using seeding for reproducibility. A success rate was calculated, indicating the percentage of articles for which a secret message was successfully extracted after modification. Next, the process was repeated with replacements of 20% up to 90% in steps of 10 percentage points. This resulted in a total of 180 000 000 processed articles for the test (10 algorithms $\times$ 9 modification percentages $\times$ 2 execution runs for short and long secret message $\times$ 1 000 000 articles). The results are presented numerically in Table 6 and graphically in Fig. 9.
Table 6: Modification Robustness
Algo. / Short Secret Message Long Secret Message
Modification 10% 20% 30% 40% 50% 60% 70% 80% 90% 10% 20% 30% 40% 50% 60% 70% 80% 90%
SNOW 80% 68% 57% 47% 37% 27% 18% 10% 3% 1% 0% 0% 0% 0% 0% 0% 0% 0%
UniSpaCh 77% 67% 57% 47% 37% 28% 18% 10% 3% 16% 11% 7% 4% 3% 1% 1% 0% 0%
AITSteg 86% 76% 66% 56% 46% 36% 26% 16% 7% 25% 17% 12% 7% 4% 2% 1% 0% 0%
Shiu et al. 4% 3% 3% 2% 2% 2% 1% 1% 0% 0% 0% 0% 0% 0% 0% 0% 0% 0%
Rizzo et al. 85% 76% 67% 59% 50% 41% 32% 22% 13% 3% 2% 2% 1% 1% 0% 0% 0% 0%
StegCloak 88% 78% 68% 58% 48% 38% 28% 18% 8% 34% 26% 18% 12% 8% 4% 2% 1% 0%
Lookalikes 11% 10% 10% 9% 8% 7% 5% 4% 2% 0% 0% 0% 0% 0% 0% 0% 0% 0%
CovertSYS 82% 72% 62% 52% 42% 32% 22% 13% 4% 32% 23% 16% 11% 6% 3% 1% 0% 0%
Shazzad-Ur-Rahman et al. 84% 74% 64% 54% 44% 34% 24% 14% 5% 4% 3% 2% 1% 1% 0% 0% 0% 0%
Innamark 86% 82% 81% 78% 75% 70% 63% 53% 34% 1% 1% 1% 0% 0% 0% 0% 0% 0%
@@
For short secret messages, our proposed Innamark technique was the most robust method in our testbed, even at high modification levels. Lookalikes and Shiu et al. had a particularly low success rate, but the other algorithms show a similar linear robustness trend. For long secret messages, all tested techniques are roughly similar, with StegCloak showing the best results because the algorithm embeds the secret message at a single position.
@@ 5.4.2 Usage Robustness
First, we analyzed a text reproduction or retyping attack [24, 33], in which users manually type the stego text in a new document. None of the algorithms could withstand this attack since they all use specific Unicode symbols that are lost on retyping.
@@
Third, and in accordance with the whitespace evaluation mentioned in Section 2.2, we analyzed the robustness of a cover text with an embedded secret message when exported to different applications and file formats. Simple copy and paste operations form “one of the most common attacks in that the malicious users copy the whole of text and paste into their own files” [33, p. 7]. Following [43], we extended the set of business-related targets to social media applications, namely, WhatsApp, Facebook Messenger, and X (formerly Twitter).
In our testbed, we applied each algorithm to some Lorem ipsum dummy text to hide a secret message in the cover text. We copied each result into the relevant application and checked whether the secret message could still be extracted after copying it back to our testbed GUI. In Table 7, “✓” indicates that the secret message could fully be extracted, whereas “✗” indicates a corrupted output. Edge cases are depicted as “(✓),” like CovertSYS in a .docx document, in which additional characters were shown in the extracted result but the original secret message could still be recognized. Only our proposed Innamark technique worked partially in the PDF format, depending on the PDF viewer used. For example, some whitespace characters were replaced with a standard U+0020 space when copying the content from Adobe Acrobat Reader, whereas they remained the same with PDF24 Reader. In such cases, “(✓)” is shown in Table 7 because the respective whitespace characters remain in the original PDF file but the robustness depends on the PDF viewer used.
Table 7: Usage Robustness

### v3

@@ 2.1 Watermarking and Steganography
Table 1: Digital watermarking definitions.
@@ 2.2 Notation and Unicode Whitespace Characters
Table 2: Whitespace evaluation based on [28].
@@
Table 3: Overview of notation, following [27].
@@
CTCT Cover text
SMSM Secret message text
@@
$CT_{SM}$ Cover text containing a hidden secret message
@@ 3.1.8 CovertSYS
Table 4: Overview of Related Methods.
@@
Innamark [19, 45] 2025 Paper, Software Replaces whitespaces with a specific set of robust, similar-looking whitespace characters.
@@ 4 Proposed Method
In this section, we present Innamark, our invisible and blind information-hiding technique for plain text. Since existing methods either lack robustness in some applications, increase the number of characters, or are recognizable by humans, they are unsuitable for the LLM use case described in Section 1. Therefore, this section presents the embedding and extraction method in Algorithm 1 and Algorithm 2 based on our nomenclature introduced in Section 2.2. It further includes a concrete example based on a Lorem ipsum dummy text and information about our implemented prototypes.
@@ 4.1 Embedding
[…]
@@
9:   $SM_{H}\leftarrow SM_{H}+a_{r+1}$ $\triangleright$ $a_{r+1}\in\mathcal{A}_{-}$
@@
16:   if $i=0$ then
17:    $CT_{SM}\leftarrow CT_{SM}+\phi$
18:    $i\leftarrow i+1$
19:   else if $0<i\leq|SM_{H}|$ then
20:    $w_{H_{i}}\leftarrow SM_{H_{i}}$ $\triangleright$ $w_{H_{i}}\in\mathcal{A}_{-}$
21:    $CT_{SM}\leftarrow CT_{SM}+w_{H_{i}}$
22:    $i\leftarrow i+1$
23:   else
24:    $i\leftarrow 0$
25:   end if
26:  else
27:   $CT_{SM}\leftarrow CT_{SM}+c$
28:  end if
29:end for
30:return  $CT_{SM}$
Algorithm 1 Embedding $Emb(CT,SM_{bytes},\theta)$
The embedding function $Emb(CT,SM_{bytes},\theta)$ begins by analyzing the configuration parameter $\theta$. This specifies the type of the secret message, which we call a InnamarkTag. It offers optional functionalities that can be enabled in any combination by the end user, namely, encryption, compression, hashing, and error-correcting codes. Depending on the user’s choice, $\theta$ defines the type used to calculate the tag. The tag has a fixed length of one byte and is returned by the $applyTag$ method with the updated secret message, depending on the user’s choice. Each bit in the tag indicates whether an option, such as compression, is enabled ($1$) or not (0). Thus, the tag describes the format of the secret message, similar to tags or headers in network packets [46]. Fig. 1 provides an overview of the InnamarkTag structure, the definition of each bit in the tag, and an example of a tag with enabled compression (second compression bit set to one) and error correction (fourth CRC32 bit set to one).
@@
An InnamarkTag starts after the separator character $\phi$, with the tag having a fixed length of one byte, followed by an optional prefix whose content depends on the tag, followed by the secret message itself.
Next, the encoding process starts transforming the secret message with the InnamarkTag and optional prefixes $SM_{bytes}$ into a sequence of whitespace characters $SM_{H}$ from our homoglyph alphabet $\mathcal{A}_{-}$. Since $|\mathcal{A}_{-}|=4$, where $|\cdot|$ denotes the cardinality, each byte of the secret message is represented by four elements of $\mathcal{A}_{-}$ because
@@
The final version of the cover text with the hidden secret message $CT_{SM}$ is created by replacing all $\delta$ successively with the elements of $SM_{H}$. If, on the one hand, the input cover text $CT$ does not have any normal whitespace $\delta$, e.g., if the algorithm has already been applied to it, it is not possible to embed the secret message. If, on the other hand, $CT$ has more $\delta$ characters than $|SM_{H}|$, the insertion process starts again until all $\delta$ are replaced to include the secret message multiple times. Thus, the resulting text does not contain standard space characters:
@@
An example of the proposed embedding algorithm is shown in Fig. 2 for the process and in Fig. 6 for input/output comparison. The secret message $SM=$ “John” is to be hidden inside the cover text $CT=$ “Lorem ipsum …” with an empty configuration parameter $\theta$ to use the default InnamarkTag without compression, hashing, or error correction. Each character of $SM$ is encoded into its byte representation $SM_{bytes}$ and transformed into the whitespace alphabet. In this case, the first letter “J” of the secret message is represented by the UTF-8 hexadecimal value 4A (U+004A), which equals the decimal value 74. Afterward, each value of $SM_{bytes}$ is encoded into $SM_{H}$ by transforming it into the alphabet $\mathcal{A}_{-}$ with a loop-based modulo operation, as described in Algorithm 1. This uses the length of the alphabet without separator characters as the divisor $d=|\mathcal{A}_{-}|=4$, as illustrated in Fig. 3. The remainder of each division operation indicates the index of the whitespace in $\mathcal{A}_{-}$ needed to build the complete sequence of whitespace characters $SM_{H}$ as a representation of the secret message $SM$. Next, each whitespace character $\delta$ of $CT$ is replaced with the corresponding space in $SM_{H}$. Since $CT$ has more whitespace characters than needed, the insertion process starts again with the separator character $\phi$, represented by the black U+2004 in Fig. 2. Afterward, it inserts the default tag and the first character “J” of the secret message a second time and stops after all whitespace characters have been replaced.
@@ 4.2 Extraction
The third part decodes the hidden secret message $SM_{H}$ into its byte representation $SM_{bytes}$. The step size for the decoding part depends on the length of the alphabet without the separator character and is defined as
@@
10: Extracted secret message bytes ($SM_{bytes}$) or error
[…]
@@
[…]
@@
[…]
@@
19:  for $y\leftarrow 0$ to d1d-1 do
20:   $a_{r+1}\leftarrow SM_{H_{i+y+1}}$ $\triangleright$ $a_{r+1}\in\mathcal{A}_{-}$
21:   $b\leftarrow b+r\cdot d^{y}$ $\triangleright$ $r\in[0,\dots,d-1]$
@@ 5.2 Capacity
We based the calculations on the approach described by Rizzo et al. [22, p. 13]: “The embedding capacity is computed as the average ratio between the number of embedded bits and the number of characters in each document.” Our dataset has an average cover text size of $\scriptstyle\sim$2514 characters. The resulting capacity ratios are depicted in Fig. 5.
@@
We assign all unbounded-capacity algorithms a value of $1.0$ to illustrate the unprescribed limit. Comparing the others, we find that Shiu et al. has the lowest embedding capacity, with $26/2514\approx 0.01$ bits/character, and that UniSpaCh has the highest of $1983/2514\approx 0.79$ bits/character. The other four bounded-capacity algorithms are close to each other and have an embedding capacity of around and below 0.1 bits/character, whereas our proposed Innamark method has a capacity of $93/2514\approx 0.04$ bits/character.
@@ 5.3 Imperceptibility
Imperceptibility or invisibility is the ability for a secret message to be concealed in a cover document without causing any visible abnormalities [21]. Therefore, this benchmark is more relevant for steganography use cases than watermarking [6, 11]. It is essential to distinguish between imperceptibility to humans and to machines, which can detect differences, for example, by using statistical metrics. Due to the varying perceptions of humans, this benchmark “is the most subjective of all the metrics” [34, p. 121]. Fig. 6 shows an illustrative example of our Innamark algorithm as a direct comparison of a plain cover text and a cover text with the embedded secret message “John” in Microsoft Office Word version 2408 using the default font “Calibri (Body)” in font size 11. To further analyze and compare the imperceptibility of all algorithms, we use four different measurement metrics in the following, namely:
@@
Figure 6: Innamark comparison example in Microsoft Word.
@@ 5.3.1 Jaro–Winkler Similarity
A standard numerical measurement used to compare the similarity between two character sequences is the Jaro–Winkler similarity, also known by the misleading name “Jaro–Winkler distance” [51]. It is often used to evaluate information-hiding techniques (see [42, 31, 27, 52, 53, 54]). The benchmark is based on the Jaro string comparator $\Phi$, for which a value of $1$ indicates that two strings $s_{1}$ and $s_{2}$ are identical and 0 indicates that the strings have no common characters [55] :
@@ 5.3.3 File Size
Humans and systems can detect a text file with a hidden secret message by its increased file size. In particular, suspicions may be raised if a file with a small amount of text has a large file size. Thus, Majeed et al. [8] argue for developing data-hiding methods that create results with minimal file sizes.
In our testbed, we compared the differences in file size between the original cover text and the text including a secret message. Although these benchmark criteria are closely related to the previous analysis of the number of characters, the results differ due to varying character storage requirements. For example, Shazzad-Ur-Rahman et al.’s [43, 44] method replaces a classical small Latin letter “g” (U+0067) with the visually similar-looking mathematical alphanumeric symbol U+1D5C0 as part of the letterlike symbols in the Unicode standard [20]. Whereas the first needs one byte of storage space in the UTF-8 encoding, the second needs four bytes. Fig. 9 illustrates the mean absolute differences based on the two runs on our 1 000 000-article dataset. The negative value for short secret messages from the Shiu et al. algorithm indicates a decreased file size. This can be attributed to replacing whitespace characters with newlines, as seen in the results for the number of characters. Our proposed Innamark technique is just in the lower third in direct comparison because it is one of the few algorithms that embeds the secret message multiple times in the cover text for increased robustness.
@@ 5.3.4 Caret Navigation
In our testbed, we manually checked various texts with secret messages. We moved the caret from the start to the end to identify any unusual behavior. In cases of hanging caret positions or identifying unexpected characters, we classified the algorithm as visible because the caret attack revealed the secret message’s position. Table 5 shows an overview of the results with a justification for each algorithm as to why it is noticeable by users or not. It was found that only three algorithms (Rizzo et al., Lookalikes, and Innamark) are invisible against caret navigations.
Table 5: Caret navigation attack evaluation.
@@ 5.4.1 Modification Robustness
Modifications to a file that may compromise the secret message, such as text insertions or deletions, are also known as tampering attacks in the information-hiding literature [33]. We have analyzed replacements, each consisting of a deletion followed by an insertion, for all implemented algorithms in batch runs for both the short and long secret messages on the 1 000 000-article dataset. For each article, we embedded the secret message and then replaced a block of 10% of the length of the original cover text with just one letter. The starting position of the replacement was randomly chosen using seeding for reproducibility. A success rate was calculated, indicating the percentage of articles for which a secret message was successfully extracted after modification. Next, the process was repeated with replacements of 20% up to 90% in steps of 10 percentage points. This resulted in a total of 180 000 000 processed articles for the test (10 algorithms $\times$ 9 modification percentages $\times$ 2 execution runs for short and long secret message $\times$ 1 000 000 articles). The results are presented in Fig. 10.
@@
For short secret messages, our proposed Innamark technique proved to be the most robust method in our testbed, even at high modification levels. Lookalikes and Shiu et al. had a particularly low success rate, but the other algorithms show a similar linear robustness trend. For long secret messages, all tested techniques are roughly similar, with StegCloak showing the best results because the algorithm embeds the secret message at a single position.
@@ 5.4.2 Usage Robustness
First, we analyzed a text reproduction or retyping attack [24, 33], in which users manually type the text with an embedded secret message in a new document. None of the algorithms could withstand this attack since they all use specific Unicode symbols that are lost on retyping.
@@
Third, and in accordance with the whitespace evaluation mentioned in Section 2.2, we analyzed the robustness of a cover text with an embedded secret message when used in different applications and file formats. Simple copy and paste operations form “one of the most common attacks in that the malicious users copy the whole of text and paste into their own files” [33, p. 7]. Following [43], we extended the set of business-related targets to social media applications, namely, WhatsApp, Facebook Messenger, and X (formerly Twitter).
In our testbed, we applied each algorithm to a Lorem ipsum dummy text to hide a secret message in the cover text. We copied each result into the relevant application and checked whether the secret message could still be extracted after copying it back to our testbed GUI. In Table 6, “✓” indicates that the secret message could fully be extracted, whereas “✗” indicates a corrupted output. Edge cases are depicted as “(✓),” like CovertSYS in a .docx document, in which additional characters were shown in the extracted result but the original secret message could still be recognized. Only our proposed Innamark technique worked partially in the PDF format, depending on the PDF viewer used. For example, some whitespace characters were replaced with a standard U+0020 space when copying the content from Adobe Acrobat Reader, whereas they remained the same with PDF24 Reader. In such cases, “(✓)” is shown in Table 6 because the respective whitespace characters remain in the original PDF file but the robustness depends on the PDF viewer used.
Table 6: Usage robustness.